OpenStudy (anonymous):

what is the height of a composite solid with a width of 3.5 ft length of 10 ft and a volume of 70 ft3

OpenStudy (wolf1728):

h x 3.5 x 10 = 70 cubic feet h = 70 / (3.5 x 10) h = 70 / 35 h = 2 feet

OpenStudy (anonymous):

thanks